OpenStudy (anonymous):

\[\int_{3}^{4} \frac{x^3 -2x^2 -4}{x^3 -2x^2}dx\]

OpenStudy (unklerhaukus):

\[\int_{3}^{4} \frac{x^3 -2x^2 -4}{x^3 -2x^2}dx\] \[=\int_{3}^{4} 1-\frac{4}{x^3 -2x^2}dx\]

OpenStudy (anonymous):

this is how far I've gotten, not sure if it's right \[\int_{3}^{4} 1+\frac{1}{x}-\frac{3}{x-2}dx\]

myininaya (myininaya):

\[\frac{-4}{x^3-2x^2}=\frac{-4}{x^2(x-2)}=\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x-2}\]

myininaya (myininaya):

By the way you can always check yourself by combining you fractions to see if you ended up with what you started out with But anyways I think you are missing x^2 on bottom somewhere

myininaya (myininaya):

Which means you broke it down wrong

OpenStudy (anonymous):

yep that's what I did...no actually \[1+\frac{Ax}{x^2}+\frac{B}{x-2}\]

myininaya (myininaya):

Yeah you are missing the over x^2 part

OpenStudy (anonymous):

where?

myininaya (myininaya):

Look what I wrote for you above Find constants A,B,and C such that above is true

OpenStudy (anonymous):

I had to do long division first because the numerator was equal to the denominator and I got \[1+\frac{4}{x^3 -2x^2}\]

OpenStudy (anonymous):

i mean the powers

OpenStudy (anonymous):

x^3

myininaya (myininaya):

Right

myininaya (myininaya):

Now you need to find And I didn't do the long division part so I don't know who is right about you or Unkle But anyways the bottom factors like this x^2(x-2) So that means you must do \[\frac{4}{x^2(x-2)}=\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x-2}\]

myininaya (myininaya):

Actually Unkle is right You don't even need long division for this

OpenStudy (anonymous):

so it wouldn't be \[\frac{Ax}{x^2}\] but instead \[\frac{A}{x}+\frac{B}{x^2}\] ?

myininaya (myininaya):

\[\frac{x^3-2x^2-4}{x^3-2x^2}=\frac{x^3-2x^2}{x^3-2x^2}+\frac{-4}{x^3-2x^2}\] \[1+\frac{-4}{x^3-2x^2}\]

OpenStudy (anonymous):

ok

myininaya (myininaya):

Yes it is what I said I'm 100 percent sure

myininaya (myininaya):

Try combining what you did, do you get what you started with The answer is no

OpenStudy (anonymous):

all right

OpenStudy (anonymous):

thanks @myininaya !

OpenStudy (anonymous):

by combining you mean...?

OpenStudy (anonymous):

just adding the terms and seeing if I get what I started with?

myininaya (myininaya):

Oh come on you know how to combine fractions :p For example to combine the following fractions (to perform the addition we need to find a common denominator) \[\frac{5}{x-2}+\frac{4}{x}\] \[\frac{5x}{x(x-2)}+\frac{4(x-2)}{x(x-2)}=\frac{5x+4(x-2)}{x(x-2)}\] Example if we wanted to do partial fractions on: Just like if we had \[\frac{4}{x(x-2)^2}=\frac{A}{x}+\frac{B}{x-2}+\frac{C}{(x-2)^2}\]